# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v3

@@ Abstract
articletype: ARTICLE PREPRINT
@@ 1 Introduction
In recent decades, remote sensing measurements within forests have been increasingly used to capture detailed forest information. Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (Bienert et al. 2007) and mobile solutions, e.g., handheld devices (Ryding et al. 2015; Bauwens et al. 2016; Marselis et al. 2016; Balenović et al. 2021), backpacks (Liang et al. 2014b, 2018; Hyyppä et al. 2020b), ground-based vehicles (Forsman, Holmgren, and Olofsson 2016; Liang et al. 2018; Bienert et al. 2018) or drones (Chisholm et al. 2013; Brede et al. 2017; Wieser et al. 2017; Hyyppä et al. 2020a). Similarly, the use of structure-from-motion (SfM) based point clouds for forest mapping has been demonstrated using different camera solutions (e.g., Liang et al. 2014a; Wallace et al. 2016; Kuželka and Surovỳ 2018; Piermattei et al. 2019). Among these methods, drone-based solutions offer several benefits, including the ability to move in areas with challenging terrain and the high potential for automatization.
@@
Hence, the use of drones in forest research has increased rapidly and offers a practical solution to collect forest data above the forest canopy (Puliti et al. 2015; Tuominen et al. 2017). Modern commercial drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems. However, their positioning systems typically rely exclusively on GNSS signals, limiting autonomous flights to open areas where these signals are reliable. In dense forests, GNSS-based localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al. 2010). Furthermore, navigating within forests requires the ability to autonomously avoid obstacles such as trees and bushes along the flight path. Consequently, drone data collection flights under the forest canopy are still primarily controlled manually by human pilots (e.g., Hyyppä et al. 2020a; Liang et al. 2019; Krisanski, Taskhiri, and Turner 2020; Prabhu et al. 2024).
@@
Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in cluttered GNSS-denied environments. Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments (e.g., Liu et al. 2022; Loquercio et al. 2021; Zhou et al. 2022). However, the proposed methods are typically validated by reporting a single successful test flight in a specific forest environment. Most studies assess the effects of varying object density only in simulators, with only a couple of methods tested through multiple test flights inside real forests.
@@
Moreover, most of the studies proposing methods for autonomous under-canopy drone trajectory planning have the scope solely on aerial robotics, and the sensor data is primarily utilized only for navigation. Only a few pioneering studies have yet explored the potential of custom robotic under-canopy drone systems in forestry applications (Liu et al. 2022; Liang et al. 2024; Cheng et al. 2024). Notably, all of these studies have utilized Ouster OS1 3D LiDAR (Ouster, San Francisco, CA, USA) as the main sensor. Whereas LiDARs have their strengths, such as 360° horizontal field of view (FOV), long sensing range, and robustness against varying illumination, they tend to be heavier and have higher power consumption than cameras (Reddy Cenkeramaddi et al. 2020). In the case of drone development, where optimizing the weight and battery consumption plays a key role, these factors lead to larger and heavier drone platforms and flight batteries that require more space between obstacles to fly safely inside the forest. Furthermore, LiDARs with suitable quality are more expensive than camera systems with built-in 3D imaging capability, leading to increased building cost. The data collected with the LiDARs is also limited to point clouds only, whereas the images collected with cameras can also be used for image-based analysis.
@@ 2.1 DBH estimation from under-canopy drone data
In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total volume of the stem alone or in combination with the three heights (Laasasenaho 1982).
@@
In the existing literature, LiDAR-based methods have been extensively studied to efficiently and accurately estimate various tree attributes, such as DBH, tree height, and stem volume. To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (Ryding et al. 2015; Bauwens et al. 2016; Cabo et al. 2018; Chen et al. 2019; Hyyppä et al. 2020b, a, 2022), though most studies have focused on relatively sparse forests with little understory vegetation.
@@
Chisholm et al. (2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually piloted drone flying under the forest canopy. The utilized custom-built drone system included a Hokuyo UTM-30LX 2D LiDAR (Hokuyo, Osaka, Japan).
@@
Hyyppä et al. (2020a) collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone in two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha. Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory. In a follow-up study, Hyyppä et al. (2020c) compared various mobile laser scanning methods, including backpack, handheld, Under-Canopy drone, and Above-Canopy drone, for forest data collection. In this study, the same manually piloted under-canopy drone system was utilized in the same forest plots as in Hyyppä et al. (2020a).
@@
Wang et al. (2021) combined data from above and under canopy flights of a manually piloted drone equipped with a Riegl miniVUX laser scanner (RIEGL Laser Measurement Systems GmbH, Horn, Austria). The study area was a boreal forest plot with an approximate tree density of 200 trees/ha.
@@
Muhojoki et al. (2024) compared the accuracy of the estimated tree attributes of several LiDAR-based mobile systems, including two commercial under-canopy drone systems, a Hovermap system (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a Deep Forestry system (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev. 5 laser scanner (Ouster, San Francisco, CA, USA). The Deep Forestry drone was flown manually, but the Hovermap drone system was able to fly autonomously in sparse forests where the distance between trees was at least 3 m. Both drone systems were used to collect point clouds under the canopy of three sparse forest plots.
@@
In addition to manually controlled drones and large autonomous commercial drone systems, recently a few pioneering studies have applied smaller custom robotic drone platforms with a high automation level for LiDAR-based under-canopy DBH estimation. Liang et al. (2024) applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha. To enable autonomous flight, a predefined global path was planned, in which the drone system made real-time adjustments during the flight according to real-time environmental conditions.
@@
Liu et al. (2022) employed a LiDAR-based semantic lidar odometry and mapping framework (SLOAM) (Chen et al. 2020) algorithm to calculate the number of trees in the explored area. Although in this study autonomous flight was demonstrated without the SLOAM algorithm and tree number calculation was only demonstrated with one manually piloted test flight, the same drone system has later been used in various follow-up studies for tree parameter estimation. Prabhu et al. (2024) extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones, as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data. However, DBH data collection flights were performed in manual flight mode. Cheng et al. (2024) used the same drone platform to collect a dataset for tree mapping based on LiDAR. In addition to data from orchards, the dataset contains the onboard LiDAR data from two forest areas. Although the drone was manually piloted in one of these forests, data collection in the other forest consisting of loblolly pine trees was carried out in autonomous flight mode. They also estimated DBH values for the LiDAR data collected during both autonomous and manual flights.
@@
During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH. Although several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al. (2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al. (2018); Piermattei et al. (2019); Xu, Shen, and Cao (2023); Gao and Kan (2022)) and above-canopy photogrammetric systems (e.g., Wallace et al. (2016); Ye, van Leeuwen, and Nyktas (2019)) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; Shimabuku, Konoshima, and Ota 2023; He et al. 2025).
@@
Kuželka and Surovỳ (2018) evaluated two manually operated drone systems, a DJI Phantom 4 Pro (DJI, Shenzhen, China) and a DJI Mavic Pro (DJI, Shenzhen, China), and various analysis methods to estimate DBH from photogrammetric point clouds. The systems were evaluated in two sparse forest plots with tree densities of 270-290 trees/ha.
@@
Krisanski, Taskhiri, and Turner (2020) performed manual flights with a DJI Phantom 4 Pro in two forest plots with stem densities of 885 trees/ha and 980 to acquire high-density SfM point cloud data and estimated the DBH of individual trees from the point clouds.
@@
Shimabuku, Konoshima, and Ota (2023) performed DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China). The drone system was able to detect obstacles in the horizontal direction, but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
@@
Recently, He et al. (2025) mounted a 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) as an additional payload to the same LiDAR-based autonomous drone platform used in Liu et al. (2022) and post-processed the photogrammetric data to 3D point clouds. However, in the study, it was not mentioned whether the flights were performed in autonomous flight mode or with manual pilot mode.
@@
The RMSEs of the DBH estimates in the relevant studies are presented in Table 1. From the studies where multiple analysis methods were evaluated, the lowest RMSEs are presented. The study by He et al. (2025) was omitted from the table, as the study reported only the median relative error (11.5 %) of the accuracy of the DBH estimation instead of the RMSE. Although some of the relevant studies also estimated tree volumes by flying the flight path several times with different altitudes, with rotating sensors, or with LiDARs with high vertical FOV, only DBH estimation accuracies are included in the table.
@@
Table 1: Accuracy of the stem diameter breast height (DBH) estimations in other studies. SfM-MVS: Structure-from-motion and multi-view stereo; ULS: Under canopy laser scanning; ALS: Above canopy laser scanning. *averaged values from three plots. The study by He et al. (2025) was omitted from this table since they reported only median errors instead of RMSE.
@@ 2.2 Autonomous flying inside forests
Even though only a few studies have yet utilized robotic under-canopy drones in DBH estimation, drone autonomy itself has been an actively researched topic in the robotics community during the past few years. Recent advances in mobile robotics and the improved computational capabilities of small embedded computers have made it realistic to build drones capable of real-time trajectory planning in cluttered GNSS-denied environments. Although technology is still in the early stages of development, some recent studies have proposed various methods that have been successfully tested in forest environments. Most importantly, the authors of these methods have also shared their source codes to make them openly available to the research community to encourage more development in the field. Although the methods utilize various sensor setups, they can be divided into LiDAR-based methods (Liu et al. 2022; Prabhu et al. 2024; Ren et al. 2022; Liu, Ren, and Zhang 2024; Ahmad et al. 2022) and camera-based methods (Campos-Macías et al. 2021; Zhou et al. 2022; Loquercio et al. 2021; Nguyen et al. 2023) based on the sensor utilized for the detection of obstacles in the flight path. Since GNSS is not a suitable method for localization under the forest canopy, these methods rely on various Simultaneous localization and mapping (SLAM) or odometry algorithms and onboard sensor data to obtain the pose estimate in a local coordinate system. Most of these methods tackle the problem of collision-free trajectory planning using 3D grid maps, but the solutions by Loquercio et al. (2021) and Nguyen et al. (2023) replaced individual mapping and trajectory planning subtasks by deep neural networks mapping the depth images directly to collision-free trajectories. In addition, the methods proposed by Zhou et al. (2022) and Ahmad et al. (2022) concentrated a swarm of drones instead of the autonomous flying of an individual drone.
However, most of these methods were validated by a single successful autonomous test flight in a specific forest environment. Only four of the articles reported multiple test flights inside real forests (Loquercio et al. 2021; Ren et al. 2022; Liu, Ren, and Zhang 2024; Prabhu et al. 2024). From these, (Loquercio et al. 2021) and (Prabhu et al. 2024) were the only methods tested at least in two different forest types. Furthermore, since the methods by (Loquercio et al. 2021), (Ren et al. 2022), and (Liu, Ren, and Zhang 2024) were all targeting high-speed flight, the test forests reported in these studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation. The autonomous trajectory planning method by Prabhu et al. (2024) was also tested only in environments where the space between the trees was larger than in typical forests, and the data-collecting flights in denser forests were manually piloted.

### v4

@@ Abstract
articletype: This is an Accepted Manuscript of an article published by Taylor & Francis in International Journal of Remote Sensing on 03 Dec 2025. The published final Version of Record is openly available at: https://doi.org/10.1080/01431161.2025.2579803.
@@ 1 Introduction
In recent decades, remote sensing measurements within forests have been increasingly used to capture detailed forest information. Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (bienert2007tree) and mobile solutions, e.g., handheld devices (ryding2015assessing; bauwens2016forest; marselis2016deriving; balenovic2021hand), backpacks (liang2014possibilities; liang2018situ; hyyppa2020accurate), ground-based vehicles (forsman2016tree; liang2018situ; bienert2018comparison) or drones (chisholm2013uav; brede2017comparing; wieser2017case; hyyppa2020under). Similarly, the use of structure-from-motion (SfM) based point clouds for forest mapping has been demonstrated using different camera solutions (e.g., liang2014use; wallace2016assessment; kuvzelka2018mapping; piermattei2019terrestrial). Among these methods, drone-based solutions offer several benefits, including the ability to move in areas with challenging terrain and the high potential for automatization.
@@
Hence, the use of drones in forest research has increased rapidly and offers a practical solution to collect forest data above the forest canopy (puliti2015forestUAS; tuominen2017hyperspectral). Modern commercial drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems. However, their positioning systems typically rely exclusively on GNSS signals, limiting autonomous flights to open areas where these signals are reliable. In dense forests, GNSS-based localization becomes unreliable due to signal blockages and multipath effects caused by trees (gnss_multipath). Furthermore, navigating within forests requires the ability to autonomously avoid obstacles such as trees and bushes along the flight path. Consequently, drone data collection flights under the forest canopy are still primarily controlled manually by human pilots (e.g., hyyppa2020under; liang2019forest; krisanski2020photodrone; prabhu2024treeFalcon).
@@
Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in cluttered GNSS-denied environments. Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments (e.g., LiuEtAl; loquercio2021learning; zhou2022swarm). However, the proposed methods are typically validated by reporting a single successful test flight in a specific forest environment. Most studies assess the effects of varying object density only in simulators, with only a couple of methods tested through multiple test flights inside real forests.
@@
Moreover, most of the studies proposing methods for autonomous under-canopy drone trajectory planning have the scope solely on aerial robotics, and the sensor data is primarily utilized only for navigation. Only a few pioneering studies have yet explored the potential of custom robotic under-canopy drone systems in forestry applications (LiuEtAl; liang2024forest; cheng2024Treescope). Notably, all of these studies have utilized Ouster OS1 3D LiDAR (Ouster, San Francisco, CA, USA) as the main sensor. Whereas LiDARs have their strengths, such as 360° horizontal field of view (FOV), long sensing range, and robustness against varying illumination, they tend to be heavier and have higher power consumption than cameras (sensors). In the case of drone development, where optimizing the weight and battery consumption plays a key role, these factors lead to larger and heavier drone platforms and flight batteries that require more space between obstacles to fly safely inside the forest. Furthermore, LiDARs with suitable quality are more expensive than camera systems with built-in 3D imaging capability, leading to increased building cost. The data collected with the LiDARs is also limited to point clouds only, whereas the images collected with cameras can also be used for image-based analysis.
@@ 2.1 DBH estimation from under-canopy drone data
In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total volume of the stem alone or in combination with the three heights (laasasenaho1982taper).
@@
In the existing literature, LiDAR-based methods have been extensively studied to efficiently and accurately estimate various tree attributes, such as DBH, tree height, and stem volume. To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (ryding2015assessing; bauwens2016forest; cabo2018comparing; chen2019applicability; hyyppa2020accurate; hyyppa2020under; hyyppa2022direct), though most studies have focused on relatively sparse forests with little understory vegetation.
@@
chisholm2013uav were the first to demonstrate DBH measurements from point cloud data collected with a manually piloted drone flying under the forest canopy. The utilized custom-built drone system included a Hokuyo UTM-30LX 2D LiDAR (Hokuyo, Osaka, Japan).
@@
hyyppa2020under collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone in two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha. Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory. In a follow-up study, hyyppa2020comparison compared various mobile laser scanning methods, including backpack, handheld, Under-Canopy drone, and Above-Canopy drone, for forest data collection. In this study, the same manually piloted under-canopy drone system was utilized in the same forest plots as in hyyppa2020under.
@@
wang2021seamless combined data from above and under canopy flights of a manually piloted drone equipped with a Riegl miniVUX laser scanner (RIEGL Laser Measurement Systems GmbH, Horn, Austria). The study area was a boreal forest plot with an approximate tree density of 200 trees/ha.
@@
muhojoki24BENCHMARKING compared the accuracy of the estimated tree attributes of several LiDAR-based mobile systems, including two commercial under-canopy drone systems, a Hovermap system (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a Deep Forestry system (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev. 5 laser scanner (Ouster, San Francisco, CA, USA). The Deep Forestry drone was flown manually, but the Hovermap drone system was able to fly autonomously in sparse forests where the distance between trees was at least 3 m. Both drone systems were used to collect point clouds under the canopy of three sparse forest plots.
@@
In addition to manually controlled drones and large autonomous commercial drone systems, recently a few pioneering studies have applied smaller custom robotic drone platforms with a high automation level for LiDAR-based under-canopy DBH estimation. liang2024forest applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha. To enable autonomous flight, a predefined global path was planned, in which the drone system made real-time adjustments during the flight according to real-time environmental conditions.
@@
LiuEtAl employed a LiDAR-based semantic lidar odometry and mapping framework (SLOAM) (chen2020sloam) algorithm to calculate the number of trees in the explored area. Although in this study autonomous flight was demonstrated without the SLOAM algorithm and tree number calculation was only demonstrated with one manually piloted test flight, the same drone system has later been used in various follow-up studies for tree parameter estimation. prabhu2024treeFalcon extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones, as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data. However, DBH data collection flights were performed in manual flight mode. cheng2024Treescope used the same drone platform to collect a dataset for tree mapping based on LiDAR. In addition to data from orchards, the dataset contains the onboard LiDAR data from two forest areas. Although the drone was manually piloted in one of these forests, data collection in the other forest consisting of loblolly pine trees was carried out in autonomous flight mode. They also estimated DBH values for the LiDAR data collected during both autonomous and manual flights.
@@
During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH. Although several studies have investigated terrestrial photogrammetric systems (e.g., liang2014use; surovy2016accuracy; mokrovs2018high; piermattei2019terrestrial; xu2023sfm_vs_ls; gao2022automaticDBH) and above-canopy photogrammetric systems (e.g., wallace2016assessment; ye2019analysing) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (kuvzelka2018mapping; krisanski2020photodrone; shimabuku2023diameter; he2025estimating360).
@@
kuvzelka2018mapping evaluated two manually operated drone systems, a DJI Phantom 4 Pro (DJI, Shenzhen, China) and a DJI Mavic Pro (DJI, Shenzhen, China), and various analysis methods to estimate DBH from photogrammetric point clouds. The systems were evaluated in two sparse forest plots with tree densities of 270-290 trees/ha.
@@
krisanski2020photodrone performed manual flights with a DJI Phantom 4 Pro in two forest plots with stem densities of 885 trees/ha and 980 to acquire high-density SfM point cloud data and estimated the DBH of individual trees from the point clouds.
@@
shimabuku2023diameter performed DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China). The drone system was able to detect obstacles in the horizontal direction, but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
@@
Recently, he2025estimating360 mounted a 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) as an additional payload to the same LiDAR-based autonomous drone platform used in LiuEtAl and post-processed the photogrammetric data to 3D point clouds. However, in the study, it was not mentioned whether the flights were performed in autonomous flight mode or with manual pilot mode.
@@
The RMSEs of the DBH estimates in the relevant studies are presented in Table 1. From the studies where multiple analysis methods were evaluated, the lowest RMSEs are presented. The study by he2025estimating360 was omitted from the table, as the study reported only the median relative error (11.5 %) of the accuracy of the DBH estimation instead of the RMSE. Although some of the relevant studies also estimated tree volumes by flying the flight path several times with different altitudes, with rotating sensors, or with LiDARs with high vertical FOV, only DBH estimation accuracies are included in the table.
@@
Table 1: Accuracy of the stem diameter breast height (DBH) estimations in other studies. SfM-MVS: Structure-from-motion and multi-view stereo; ULS: Under canopy laser scanning; ALS: Above canopy laser scanning. *averaged values from three plots. The study by he2025estimating360 was omitted from this table since they reported only median errors instead of RMSE.
@@ 2.2 Autonomous flying inside forests
Even though only a few studies have yet utilized robotic under-canopy drones in DBH estimation, drone autonomy itself has been an actively researched topic in the robotics community during the past few years. Recent advances in mobile robotics and the improved computational capabilities of small embedded computers have made it realistic to build drones capable of real-time trajectory planning in cluttered GNSS-denied environments. Although technology is still in the early stages of development, some recent studies have proposed various methods that have been successfully tested in forest environments. Most importantly, the authors of these methods have also shared their source codes to make them openly available to the research community to encourage more development in the field. Although the methods utilize various sensor setups, they can be divided into LiDAR-based methods (LiuEtAl; prabhu2024treeFalcon; ren2022BubblePlanner; LiuIPC2024; Ahmad2022Bioinspired) and camera-based methods (CamposEtAl; zhou2022swarm; loquercio2021learning; nguyen2023Unceratainity) based on the sensor utilized for the detection of obstacles in the flight path. Since GNSS is not a suitable method for localization under the forest canopy, these methods rely on various Simultaneous localization and mapping (SLAM) or odometry algorithms and onboard sensor data to obtain the pose estimate in a local coordinate system. Most of these methods tackle the problem of collision-free trajectory planning using 3D grid maps, but the solutions by loquercio2021learning and nguyen2023Unceratainity replaced individual mapping and trajectory planning subtasks by deep neural networks mapping the depth images directly to collision-free trajectories. In addition, the methods proposed by zhou2022swarm and Ahmad2022Bioinspired concentrated a swarm of drones instead of the autonomous flying of an individual drone.
However, most of these methods were validated by a single successful autonomous test flight in a specific forest environment. Only four of the articles reported multiple test flights inside real forests (loquercio2021learning; ren2022BubblePlanner; LiuIPC2024; prabhu2024treeFalcon). From these, (loquercio2021learning) and (prabhu2024treeFalcon) were the only methods tested at least in two different forest types. Furthermore, since the methods by (loquercio2021learning), (ren2022BubblePlanner), and (LiuIPC2024) were all targeting high-speed flight, the test forests reported in these studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation. The autonomous trajectory planning method by prabhu2024treeFalcon was also tested only in environments where the space between the trees was larger than in typical forests, and the data-collecting flights in denser forests were manually piloted.